Please, help Determine the molar solubility of Mg(OH)2 in a solution buffered at a pH of 7. KSP=1.2*10^-11
Answer #1

use:

pH = -log [H+]

7 = -log [H+]

[H+] = 1*10^-7 M

use:

[OH-] = Kw/[H+]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[OH-] = (1.0*10^-14)/[H+]

[OH-] = (1.0*10^-14)/(1*10^-7)

[OH-] = 1*10^-7 M

At equilibrium:

Mg(OH)2 <----> Mg2+ + 2 OH-

   s 1*10^-7

Ksp = [Mg2+][OH-]^2

1.2*10^-11=(s)*(1*10^-7)^2

1.2*10^-11= (s) * 10^-14

s = 1.2*10^3 M

Answer: 1.2*10^3 M